# Clustering by Denoising: Latent plug-and-play diffusion for single-cell data

### v1

@@ 1 Introduction
Single-cell RNA sequencing (scRNA-seq) has revolutionized biomedical research by enabling high-resolution profiling of cellular heterogeneity (Park et al., 2020; Miragaia et al., 2019), with large-scale initiatives like the Human Cell Atlas providing foundational references for cell type annotation  (Regev et al., 2017; Lindeboom et al., 2021; Elmentaite et al., 2022; Stuart et al., 2019; Lopez et al., 2018). However, since cells are sequenced without predefined labels, accurate cell type identification must be derived entirely from unsupervised analysis of noisy high-dimensional data. The canonical approach involves reducing dimensionality (e.g., via PCA) followed by clustering and manual annotation based on marker genes—an iterative and subjective process (Kiselev et al., 2019; Stuart et al., 2019). This paradigm suffers from three major limitations: 1) scRNA-seq data is inherently noisy due to technical artifacts (e.g., varying capture efficiency, Kharchenko et al. 2014) and biological stochasticity (Wagner et al., 2016), which clustering algorithms amplify, often producing unreliable labels; 2) Reproducibility is hampered by batch effects and procedural differences across labs (Shaham et al., 2017), making consistent clustering across datasets difficult; and 3) Population-specific and technology-induced shifts in data distribution reduce generalizability.
We frame single-cell denoising as an inverse problem: recovering clean gene expression from noisy measurements without imposing strong generative assumptions. This follows the plug-and-play (PnP) paradigm (Venkatakrishnan et al., 2013a; Zhang et al., 2021; Chan et al., 2016; Ryu et al., 2019), which integrates powerful denoising priors with measurement models. Mainstream PnP diffusion frameworks (Zhu et al., 2023; Go et al., 2023; Wu et al., 2024; Xu & Chi, 2024) enables this through combining likelihoods via iterative refinement (e.g., Gibbs sampling), where each denoising step is followed by controlled noise reintroduction to enforce data consistency. In principle, this enables denoising beyond test data quality by transferring patterns from high-signal reference data (e.g., SMART-seq2, Picelli et al. 2013; 2014) to noisier technologies (e.g., droplet-based scRNA-seq, Klein et al. 2015; Macosko et al. 2015). However, directly applying image-based PnP frameworks to single-cell data is challenging. Unlike images where pixel noise is largely independent, gene expression data exhibits intrinsic low-rank structure and complex correlations. Moreover, denoising must preserve relational structure between cells for accurate clustering and annotation. Standard dimensionality reduction (e.g., PCA) can collapse distinct cell types, making it impossible to guide denoising accurately.
@@
Our framework diverges from existing Bayesian approaches for single-cell analysis by removing their need for restrictive generative modeling. While variational autoencoders (VAEs) (Lopez et al., 2018; Gayoso et al., 2019; Grønbech et al., 2020) are difficult to train and rely on strong likelihood assumptions, more recent approximate message passing methods with empirical Bayes denoisers (Zhong et al., 2022; Nandy & Ma, 2024) still require parametric noise modeling, operate purely in the latent space, and scale poorly to high-dimensional latent spaces. In contrast, our approach requires no explicit generative model or pre-processing for noise structure, instead learning it directly from data. By combining the adaptability of likelihood-free diffusion with the structure-aware refinement of Gibbs-based input-space guidance, we enhance cluster separation even under high noise or distribution shift, without depending on restrictive architectural or noise-modeling priors.
Our framework offers three key advantages over existing single-cell clustering and denoising methods, providing a principled, robust, and reproducible pathway for automated cell type annotation:
@@
Additional Related Work on Diffusion in Single-Cell Data   Diffusion models provide flexible, trainable priors that accommodate complex noise structures (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021b; a). While initial single-cell applications have focused on transcriptome generation and data imputation (Luo et al., 2024; Wang et al., 2025; 2024; Zhang et al., 2025), their potential for learning denoised low-dimensional embeddings, critical for reference atlas construction and robust label transfer, remains largely unexplored beyond restrictive generative modelling assumptions. Our work investigates this gap, using unsupervised clustering and subsequent biological validation as the evaluation framework.
@@ 2 Single-Cell Atlas Construction via Posterior Sampling
We consider two single-cell RNA-seq datasets: (1) a reference dataset $\mathcal{D}^{{(r)}}=\left\{X^{{(r)}}_{1},\ldots,X^{{(r)}}_{m}\right\}$ used to train a diffusion prior that captures the underlying biological manifold, and (2) a target dataset $\mathcal{D}^{{(t)}}=\left\{X^{{(t)}}_{1},\ldots,X^{{(t)}}_{n}\right\}$ to which we apply denoising using the learned prior. Both datasets consist of raw UMI counts transformed into expression profiles via standard preprocessing pipelines (e.g., quality control, variance stabilization, etc.). Our framework is agnostic to specific preprocessing choices and accommodates diverse noise structures.
Data Generating Process   We assume both datasets follow a low-rank factor model (Peng et al., 2021; Weine et al., 2024; Zhong et al., 2022; Nandy & Ma, 2024; Argelaguet et al., 2020). Let $X_{i}\in\mathbb{R}^{d}$:
@@
where $V\in\mathbb{R}^{d\times k}$ is a factor loading matrix that spans the transcriptional space; $U_{i}\in\mathbb{R}^{k}\sim P_{\text{prior}}$ encodes the low-dimensional biological cell signal drawn i.i.d. from an unknown distribution $P_{\text{prior}}$; and $\varepsilon_{i}$ is an independent noise vector, reflecting stochastic variation in the measurement process or biological stochasticity. For specific datasets, we have
@@
$X^{(t)}$ $=U^{(t)}\left(V^{(t)}\right)^{\top}+\varepsilon^{(t)},\quad U^{(t)}\in\mathbb{R}^{n\times k},\quad\varepsilon^{(t)}\in\mathbb{R}^{n\times d}.$ (3)
We assume conserved latent biological structures across datasets: the prior distribution $P_{\text{prior}}$ (shared cell state manifold) and factor loading matrix $V$ (transcriptional programs) are identical for both reference and target datasets. However, we allow dataset-specific noise characteristics $\varepsilon^{(t)}\nsim\varepsilon^{(r)}$ to accommodate different measurement technologies and batch effects, with the expectation that our method performs best when the reference dataset $\mathcal{D}^{(r)}$ has comparable or lower noise levels than the target dataset $\mathcal{D}^{(t)}$.
Given these shared structures, we drop superscripts on $U$ and $V$ entirely, as they represent the same biological quantities across datasets. We retain the $(r)$ and $(t)$ superscripts only for dataset-specific terms: $\mathcal{D}^{(r)}$, $\mathcal{D}^{(t)}$, $X^{(r)}$, $X^{(t)}$, $\varepsilon^{(r)}$, and $\varepsilon^{(t)}$. To further simplify notation, we henceforth drop the superscripts with the convention that all quantities refer to the target dataset $\mathcal{D}^{(t)}$ unless explicitly marked with $(r)$ for the reference dataset $\mathcal{D}^{(r)}$.
Posterior Sampling for Denoised Embeddings   Our goal is to construct a denoised atlas of $\mathcal{D}^{{(t)}}$ by computing posterior embeddings $\mathbb{E}[U_{i}\mid X_{i}]$, which we frame as sampling from the posterior:
@@
where $f(\cdot)$ denotes the likelihood associated with the observation model, and $P_{\text{prior}}$ represents the population prior learned via diffusion on $\mathcal{D}^{{(r)}}$. The main challenge lies in the likelihood term $f(X-UV^{\top}\mid U)$, which encodes the relationship between latent embeddings and observed expressions through the factor model in Eq. (1). Traditional approaches address this difficulty by imposing restrictive assumptions: conjugate Gaussian priors with Gaussian likelihoods (Gelman et al., 2013) enable tractable inference but fail to capture complex biological distributions, while Metropolis-adjusted Langevin algorithms (Roberts & Rosenthal, 1998; Durmus et al., 2018) handle non-Gaussian likelihoods but struggle with implicitly defined diffusion priors.
PnP Framework with Auxiliary Variables   By leveraging the PnP diffusion framework, we overcome this challenge using a split Gibbs sampling approach (Xu & Chi, 2024) that introduces auxiliary variables $Z_{i}$ to decouple the likelihood from the diffusion prior. This is achieved by first replacing $U_{i}$ with $Z_{i}$ in the likelihood generating stage, and then enforce consistency between $U_{i}$ and $Z_{i}$ through a Gaussian penalty, which leads to following augmented joint distribution:
@@
    Unsupervised Clustering with Post-Hoc Evaluation: Denoised embeddings $\{{\widehat{U}_{i}}\}_{i=1}^{n}$ are clustered without using label information to obtain $K$ clusters. The quality of these clusters is evaluated by comparing them to the known cell type labels $L_{i}\in{1,\ldots,C}$ using metrics like adjusted rand index (ARI) (Hubert & Arabie, 1985), average silhouette score (Rousseeuw, 1987), and cell-type Locally Invariant Simpson Index (Korsunsky et al., 2019), which quantify the agreement between the data-driven clusters and biological annotations.
@@ 3 Method
This approach preserves the flexibility of diffusion priors while maintaining tractable posterior inference, providing a scalable and uncertainty-aware framework for single-cell atlas construction.
@@ 4 Evaluation on synthetic data
Evaluation   We assess the cluster separation of DICE-derived embeddings against a PCA baseline that projects expression profiles to the same latent dimension used for clustering. Qualitatively, we visualize UMAPs (McInnes et al., 2018) (Fig. 1); quantitatively, we report the average cosine silhouette score (Rousseeuw, 1987), computed using the true cluster labels of the test data (Table 1). Because ground-truth labels are available in this benchmark, both readouts are directly comparable. Across all four settings, DICE yields clearer separation in UMAP and higher silhouette scores than the PCA baseline, indicating more faithful recovery of the underlying classes.
@@
Setup PCA DICE
1 0.25 0.37
2 0.24 0.36
3 0.22 0.34
4 0.22 0.28
Table 1: Average silhouette scores (higher is better) for two methods across four settings.
@@ 5.1 Analysis of the CITE-seq dataset
Preprocessing and denoising   We applied standard QC, removing cells with <200<200 genes, genes expressed in <3<3 cells, and cells with $15\%$ mitochondrial counts. Using library size factors computed from the dataset, we normalized counts to $10^{4}$ per cell, followed by $\log(1+x)$ transformation to stabilize variance. We selected the top 3,000 highly variable genes (Seurat v3 criterion), scaled to unit variance with values capped at 10. Batch effects from sequencing lanes were corrected with Harmony (Korsunsky et al., 2019), yielding a matrix $X\in\mathbb{R}^{10000\times 3000}$. We split into 9,000 training and 1,000 held-out test cells with preserved label proportions. Similar to Setup 1, we test DICE’s ability in denoising beyond the quality of the training dataset in this setting. The training set was used to fit the diffusion model, and the test set was reserved for evaluation.
@@
Evaluation   We compare UMAP (McInnes et al., 2020) visualizations of the 1,000 held-out test cells denoised by DICE against their PCA projections without denoising (Figure 3). Denoised embeddings show clearer segregation of immune subtypes, with marked improvement in separating sub lineages of CD4 and CD8 T cells. The DICE atlas also performs better in seggregating the MAIT cells from the other T cells. These populations are notoriously difficult to resolve from RNA alone (Hao et al., 2021), a challenge that originally motivated multimodal approaches combining RNA and protein. In the unimodal setting considered here, however, popular toolkits such as Seurat rely directly on PCA embeddings. Our results indicate that denoising PCA embeddings with priors learned from high-signal training data substantially improves cluster separation, supporting more reliable annotation of held-out cells. Beyond visual inspection, we also quantify clustering quality using the Adjusted Rand Index (ARI) (Hubert & Arabie, 1985) computed after Leiden clustering on a 20-nearest-neighbor graph built from the embeddings, comparing recovered clusters to the true cell labels provided in the metadata. The PCA-based embeddings achieve an ARI of $0.7165$, whereas the DICE embeddings achieve $0.7563$, illustrating the effectiveness of our approach in better segregating distinct cell populations.
Figure 4: UMAP of 1,000 randomly sampled cells from the Polioudakis et al. (2019) dataset. Left: PCA embeddings in a 15-dimensional latent space using $\widehat{V}$ from the training set. Right: embeddings after denoising with DICE using a diffusion model trained on the Nowakowski et al. (2017) dataset.

### v2

@@ 1 Introduction
Single-cell RNA sequencing (scRNA-seq) has revolutionized biomedical research by enabling high-resolution profiling of cellular heterogeneity (Park et al., 2020; Miragaia et al., 2019), with large-scale initiatives like the Human Cell Atlas providing foundational references for cell type annotation  (Regev et al., 2017; Lindeboom et al., 2021; Elmentaite et al., 2022; Stuart et al., 2019; Lopez et al., 2018a). However, since cells are sequenced without predefined labels, accurate cell type identification must be derived entirely from unsupervised analysis of noisy high-dimensional data. The canonical approach involves reducing dimensionality (e.g., via PCA) followed by clustering and manual annotation based on marker genes—an iterative and subjective process (Kiselev et al., 2019; Stuart et al., 2019). However, this entire paradigm is often undermined by the inherent noise in scRNA-seq data, which arises from technical artifacts like varying capture efficiency (Kharchenko et al., 2014) and biological stochasticity (Wagner et al., 2016), which are amplified by standard clustering algorithms and lead to unreliable labels.
We frame single-cell denoising as an inverse problem: recovering clean gene expression from noisy measurements without imposing strong generative assumptions. Our method leverages the high-quality signal from reference datasets (e.g., SMART-seq2, Picelli et al. 2013; 2014) to learn a powerful denoising prior, which is then applied to enhance noisier target data (e.g., droplet-based scRNA-seq, Klein et al. 2015; Macosko et al. 2015) for improved clustering. We operationalize this under the plug-and-play (PnP) paradigm (Venkatakrishnan et al., 2013a; Zhang et al., 2021; Chan et al., 2016; Ryu et al., 2019), which integrates powerful denoising priors with measurement models. Mainstream PnP diffusion frameworks (Zhu et al., 2023; Go et al., 2023; Wu et al., 2024; Coeurdoux et al., 2024; Xu & Chi, 2024) enables this through combining likelihoods via iterative refinement (e.g., Gibbs sampling), where each denoising step is followed by controlled noise reintroduction to enforce data consistency. This principled approach enables us to denoise beyond test data quality by transferring patterns from high-signal reference data to noisier technologies, even under mild distribution shifts between the reference and target distribution. However, directly applying image-based PnP frameworks to single-cell data is challenging. Unlike images where pixel noise is largely independent, gene expression data exhibits intrinsic low-rank structure and complex correlations. Moreover, denoising must preserve relational structure between cells for accurate clustering and annotation. Standard dimensionality reduction (e.g., PCA) can collapse distinct cell types, making it impossible to guide denoising accurately.
@@
Our framework diverges from existing Bayesian approaches for single-cell analysis by removing their need for restrictive generative modeling. While variational autoencoders (VAEs) (Lopez et al., 2018a; Gayoso et al., 2019; Grønbech et al., 2020) are difficult to train and rely on strong likelihood assumptions, more recent approximate message passing methods with empirical Bayes denoisers (Zhong et al., 2022; Nandy & Ma, 2024) still require parametric noise modeling, operate purely in the latent space, and scale poorly to high-dimensional latent spaces. In contrast, our approach requires no explicit generative model or pre-processing for noise structure, instead learning it directly from data.
By combining the adaptability of likelihood-free diffusion with the structure-aware refinement of Gibbs-based input-space guidance, we enhance cluster separation in the denoised low-dimensional embeddings of the target data, even when the biological signal is weak or the target distribution moderately differs from the reference.
Our diffusion based denoising framework offers three key advantages over existing single-cell denoising methods, and therefore provides a principled, robust, and reproducible pathway for diverse downstream tasks like automated cell type annotation using the denoised embeddings.
@@
1.1 Additional Related Work
Diffusion in Single-Cell Data   Diffusion models provide flexible, trainable priors that accommodate complex noise structures (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021b; a). While initial single-cell applications have focused on transcriptome generation and data imputation (Luo et al., 2024; Wang et al., 2025; 2024; Zhang et al., 2025), their potential for learning denoised low-dimensional embeddings, critical for reference atlas construction and robust label transfer, remains largely unexplored beyond restrictive generative modelling assumptions. Our work investigates this gap, using unsupervised clustering and subsequent biological validation as the evaluation framework.
Single-Cell Preprocessing Analyzing single-cell data is challenging because raw UMI counts vary across labs, technologies, and populations (Shaham et al., 2017). Several pipelines have been proposed aim to mitigate these effects, most prominently the ubiquitous Seurat workflow (Stuart et al., 2019) applying quality control, and normalization. In datasets with pronounced batch effects, such batch-induced variations are typically corrected using techniques like Harmony (Korsunsky et al., 2019), and the batch corrected expressions are used to produce a linear embedding for downstream tasks such as clustering or label transfer. An alternative direction incorporates these steps into deep generative models that operate directly on raw counts, such as scVI (Lopez et al., 2018b), which learns latent representations end-to-end.
Instead of creating embeddings for all cells, a complementary line of work aggregates similar cells into metacells (e.g., MetaQ (Li et al., 2025)), which are then used for downstream analyses. In contrast, our method produces embeddings for all individual cells, rather than summarizing them.
@@ 2 Single-Cell Atlas Construction via Posterior Sampling
This work considers two single-cell RNA-seq datasets: (1) a reference dataset $\mathcal{D}^{{(r)}}=\left\{X^{{(r)}}_{1},\ldots,X^{{(r)}}_{m}\right\}$ used to train a diffusion prior that captures the underlying biological manifold, and (2) a target dataset $\mathcal{D}^{{(t)}}=\left\{X^{{(t)}}_{1},\ldots,X^{{(t)}}_{n}\right\}$ for denoising. Our framework is designed for single-cell RNA-seq datasets after standard preprocessing steps, such as quality control to filter low-quality cells, library-size normalization, and a $\log 1p$ transform which renders the residual non-biological variation approximately Gaussian. Within this regime, our method does not rely on stronger assumptions about upstream preprocessing choices. The goal is to construct a single-cell atlas—a reference map of cellular states— learned via denoised low-dimensional embeddings that capture underlying biological structure while removing technical variation arising from non-biological sources.
Data Generating Process   Both datasets are modeled using a low-rank factor model (Peng et al., 2021; Weine et al., 2024; Zhong et al., 2022; Nandy & Ma, 2024; Argelaguet et al., 2020). Let $X_{i}\in\mathbb{R}^{d}$ represent the gene expression profile for cell $i$:
@@
where $V\in\mathbb{R}^{d\times k}$ is a factor loading matrix that spans the transcriptional space; $U_{i}\in\mathbb{R}^{k}\sim P_{\text{prior}}$ encodes the low-dimensional biological cell signal drawn i.i.d. from an unknown distribution $P_{\text{prior}}$; and $\varepsilon_{i}$ is an independent noise vector, reflecting stochastic variation in the measurement process or biological stochasticity.
Our method is built on a generative model that projects both reference and target datasets into a shared latent space. This is formalized as:
@@
$X^{(t)}$ $=U^{(t)}\left(V^{(r)}\right)^{\top}+\varepsilon^{(t)},\quad U^{(t)}\in\mathbb{R}^{n\times k},\quad\varepsilon^{(t)}\in\mathbb{R}^{n\times d}.$ (3)
The critical modeling choice is the shared factor loading matrix $V$, learned from the reference dataset. This allows the target data to be projected into the same latent space as the reference, enabling knowledge transfer. This approach explicitly accommodates different measurement technologies through distinct noise models for each dataset ($\varepsilon^{(t)}\nsim\varepsilon^{(r)}$), with the expectation that our method performs best when the reference dataset $\mathcal{D}^{(r)}$ has comparable or lower noise levels than the target dataset $\mathcal{D}^{(t)}$.
Given this shared structure, we drop superscripts on $V$ entirely. We retain the $(r)$ and $(t)$ superscripts only for dataset-specific terms: $\mathcal{D}^{(r)}$, $\mathcal{D}^{(t)}$, $X^{(r)}$, $X^{(t)}$, $U^{(r)}$, $U^{(t)}$, $\varepsilon^{(r)}$, and $\varepsilon^{(t)}$. To further simplify notation, we henceforth drop the superscripts with the convention that all quantities refer to the target dataset $\mathcal{D}^{(t)}$ unless explicitly marked with $(r)$ for the reference dataset $\mathcal{D}^{(r)}$.
Posterior Sampling for Denoised Embeddings   Our goal is to construct a denoised atlas of $\mathcal{D}^{{(t)}}$ by computing posterior embeddings $\mathbb{E}[U_{i}\mid X_{i}]$, framed as sampling from the posterior:
@@
where $f(\cdot)$ denotes the likelihood associated with the observation model, and $P_{\text{prior}}$ represents the population prior learned via diffusion on $\mathcal{D}^{{(r)}}$. The main challenge in sampling arises from the joint influence of likelihood and prior. Efficiently sampling from a posterior that couples these two components is non-trivial, as it requires balancing the local reconstruction constraints encoded by the likelihood with the global manifold structure enforced by the prior. Traditional approaches address this difficulty by imposing restrictive assumptions: conjugate Gaussian priors with Gaussian likelihoods (Gelman et al., 2013) enable tractable inference but fail to capture complex biological distributions, while Metropolis-adjusted Langevin algorithms (Roberts & Rosenthal, 1998; Durmus et al., 2018) handle non-Gaussian likelihoods but struggle with implicitly defined diffusion priors.
Remark 2.1 (Design Principle: Reference-Guided Denoising).
A core design principle of this method is to leverage the high-quality reference dataset to learn a biologically meaningful prior, $P_{\text{prior}}(U)$, which captures the latent manifold of plausible cell states. This prior is not used to force the target dataset into an identical configuration, but to provide a structural guide for denoising. The shared projection matrix $V$ enables this by projecting both datasets into a common latent space, allowing the diffusion prior learned from $\mathcal{D}^{(r)}$ to provide a structural guide for denoising the target data. By projecting the noisier target data into this well-defined space, our diffusion prior can effectively separate biological signal from dataset-specific noise while accommodating different measurement technologies through distinct noise models.
PnP Framework with Auxiliary Variables   By leveraging the PnP diffusion framework, we overcome this challenge using a split Gibbs sampling approach (Vono et al., 2019; Xu & Chi, 2024) that introduces auxiliary variables $Z_{i}$ to decouple the likelihood from the diffusion prior. This is achieved by first replacing $U_{i}$ with $Z_{i}$ in the likelihood generating stage, and then enforce consistency between $U_{i}$ and $Z_{i}$ through a Gaussian penalty, which leads to following augmented joint distribution:
@@
    Unsupervised Clustering with Post-Hoc Evaluation: Denoised embeddings $\{{\widehat{U}_{i}}\}_{i=1}^{n}$ are clustered without using label information to obtain $K$ clusters. The quality of these clusters is evaluated by comparing them to the known cell type labels $L_{i}\in{1,\ldots,C}$ using the metrics adjusted rand index (ARI) (Hubert & Arabie, 1985), average silhouette score (Rousseeuw, 1987), normalized mutual information (NMI) (Strehl & Ghosh, 2002), and cell type locally invariant Simpson’s index (cLISI) Korsunsky et al. (2019) [^1] For detailed description of these metrics, see Section D., which quantify the agreement between the data-driven clusters and biological annotations.
@@ 3 Method
An illustration of the denoising can be found in Figure A.1. This approach preserves the flexibility of diffusion priors while maintaining tractable posterior inference, providing a scalable and uncertainty-aware framework for single-cell atlas construction.
@@ 4 Evaluation on synthetic data
Evaluation   We assess the cluster separation of DICE-derived embeddings against a PCA baseline that projects expression profiles to the same latent dimension used for clustering. Qualitatively, we visualize UMAPs (McInnes et al., 2018) (Fig. 1); quantitatively, we report the average cosine silhouette score (Rousseeuw, 1987), and cLISI (Korsunsky et al., 2019) computed using the true cluster labels of the test data (Table 1). Because ground-truth labels are available in this benchmark, both metrics are directly comparable. Across all four settings, DICE yields clearer separation in UMAP and better scores than the PCA baseline, indicating more faithful recovery of the underlying classes.
@@
Silhouette cLISI
Setup PCA DICE PCA DICE
1 0.25 0.37 1.27 1.17
2 0.24 0.36 1.27 1.17
3 0.22 0.34 1.32 1.18
4 0.22 0.28 1.35 1.27
Table 1: Comparsion of average silhouette scores (higher is better) and cLISI scores (lower is better) calculated on the ground-truth cluster labels across four synthetic setups.
@@ 5 Evaluation on single-cell data
For CITE-seq, training takes about 36 minutes and inference takes about 12 minutes on a setup accelerated by an NVIDIA RTX PRO 6000 GPU. We provide more details on our setup in Appendix C.
@@ 5.1 Analysis of the CITE-seq dataset
Preprocessing and denoising   We applied standard QC and pre-processing to the raw UMI counts data (see details in Section H), and considered the top 3,000 highly variable genes (Seurat v3 criterion) to construct the data matrix $X\in\mathbb{R}^{10{,}000\times 3{,}000}$. We split into 9,000 training and 1,000 held-out test cells with preserved label proportions. Similar to Setup 1, we test DICE’s ability in denoising beyond the quality of the training dataset in this setting. The training set was used to fit the diffusion model, and the test set was reserved for evaluation.
@@
Evaluation   We compare UMAP (McInnes et al., 2020) visualizations of the 1,000 held-out test cells denoised by DICE against their PCA projections without denoising (Figure 3). Denoised embeddings show clearer segregation of immune subtypes, with marked improvement in separating sub lineages of CD4 and CD8 T cells. The DICE atlas also performs better in seggregating the MAIT cells from the other T cells. These populations are notoriously difficult to resolve from RNA alone (Hao et al., 2021), a challenge that originally motivated multimodal approaches combining RNA and protein. In the unimodal setting considered here, however, popular toolkits such as Seurat rely directly on PCA embeddings. Our results indicate that denoising PCA embeddings with priors learned from high-signal training data substantially improves cluster separation, supporting more reliable annotation of held-out cells. We additionally benchmarked the DICE embeddings against several widely used denoising procedures for RNA-seq analysis—MAGIC (van Dijk et al., 2018a), ALRA (Linderman et al., 2022a), $k$-nearest-neighbor smoothing (used in Seurat v3 (Stuart et al., 2019)), and non-negative matrix factorization (NMF)–based embeddings [^3] See Section I for details on the benchmarking procedures and their implementation., using the clustering metrics ARI, NMI, cLISI, and V-measure described in Section D. The results, summarized in Table 2, show that DICE consistently outperforms these popular denoising pipelines across most clustering metrics.
CITE-seq Human Neo-Cortex
Method ARI cLISI NMI V Measure ARI cLISI NMI V Measure
DICE 0.805 1.344 0.740 0.813 0.393 1.406 0.553 0.559
PCA 0.745 1.334 0.689 0.787 0.347 1.496 0.496 0.525
ALRA 0.604 1.361 0.713 0.775 0.310 1.526 0.474 0.512
kNN (5 nearest neighbors) 0.713 1.334 0.647 0.756 0.275 1.837 0.349 0.405
kNN (10 nearest neighbors) 0.705 1.385 0.663 0.757 0.285 1.750 0.440 0.451
kNN (15 nearest neighbors) 0.735 1.371 0.683 0.766 0.268 1.701 0.442 0.450
MAGIC 0.674 1.397 0.648 0.757 0.317 1.451 0.502 0.509
NMF 0.448 2.332 0.430 0.556 0.209 2.237 0.220 0.291
scVI (10 latent factors) 0.641 1.361 0.595 0.715
Table 2: Comparison of clustering metrics across single-cell datasets.
Figure 4: UMAP of 15,126 cells from the Polioudakis et al. (2019) dataset. Left: PCA embeddings in a 15-dimensional latent space using $\widehat{V}$ from the training set. Right: embeddings after denoising with DICE using a diffusion model trained on the Nowakowski et al. (2017) dataset.